OpenStudy (anonymous):

simplify completely 4x^3 + 5x^2 + 3x over x?

OpenStudy (anonymous):

Take x as common

OpenStudy (callisto):

( 4x^3 + 5x^2 + 3x)/x = x(4x^2+5x+3) /x = 4x^2 +5x +3